Mary is 43. Her son is 7. How many years will he be 1/3 Mary's age?
Answer by ptaylor(2198)

Let x= number of years until Mary's son is 1/3 Mary's age
So, our equation to solve then is:
(7+x)/(43+x)=1/3 cross multiply or multiply each side by 3(43+x)
3(7+x)=43+x
21+3x=43+x subtract x and also 21 from each side
21-21+3x-x=43-21+x-x collect like terms
2x=22
x=11 yrs---number of years until Mary's son is 1/3 Mary's age
CK
(7+11)/(43+11)=1/3
18/54=1/3
1/3=1/3
Hope this helps--ptaylor